# Counterfactual Fairness through Transforming Data Orthogonal to Bias

### v1

@@ 3.3. Orthogonal to bias
$\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{k},$
where the last constraint requires $\mathbf{U}$ to be orthonormal matrix, which helps prevent degeneracy, such as basis vectors becoming identically zero or encountering solutions with double multiplicity.
As shown by (Aliverti et al., 2021), the original problem (5) can be reformulated using Lagrange multipliers as follows:
(6) $\arg\min_{\mathbf{S},\mathbf{U}}~{}\frac{1}{n}\sum_{i=1}^{n}\left\lVert\textbf{a}_{i}-\sum_{j=1}^{k}s_{ij}\textbf{u}_{j}^{\top}\right\rVert^{2}+\frac{2}{n}\sum_{j=1}^{k}\lambda_{j}\sum_{i=1}^{n}s_{ij}B_{i}.$
The factor $2/n$ is introduced to simplify the expression for the optimal solutions. It ensures that the first-order condition for (6) with respect to $s_{ij}$ involves a common factor of $2/n$, which can then be canceled out during computations. Let $\mathbf{S}^{*}$ and $\mathbf{U}^{*}$ denote the optimal solutions of $\mathbf{S}$ and $\mathbf{U}$, respectively. As demonstrated in Appendix A, the reformulated equation (6) yields a closed-form solution:
(7) $\begin{split}\lambda^{*}_{j}&=\frac{\left\langle\mathbf{A}\mathbf{u}^{*\top}_{j},\mathbf{B}\right\rangle}{\langle\mathbf{B},\mathbf{B}\rangle}\\ \mathbf{U}^{*}&=\left[\textbf{u}^{*}_{1},\dots,\textbf{u}^{*}_{k}\right]\\ \mathbf{S}^{*}&=\{s^{*}_{ij}\},\end{split}$
where $\mathbf{U}^{*}$ consists of the first $k$ right singular vectors of $\mathbf{A}$ and $s_{ij}^{*}=\mathbf{a}_{i}\mathbf{u}^{*\top}_{j}-\lambda^{*}_{j}B_{i}$. Detailed derivations of the closed-form solution (7) can be found in Appendix A. The processed non-sensitive data matrix is therefore $\widetilde{\mathbf{A}}=\mathbf{S}^{*}\mathbf{U^{*}}^{\top}$.
It is noteworthy that when the correlation between the sensitive variable and the observed data is minimal, $\widetilde{\mathbf{A}}$ yields a reconstruction error similar to that of the standard Singular Value Decomposition (SVD). It is because the additional reconstruction error of $\widetilde{\mathbf{A}}$ relative to SVD with the same rank is proportional to the collinearity between the subspace spanned by $\mathbf{B}$ and the left singular vectors of the data $\mathbf{A}$ (Aliverti et al., 2021). The exact expression for OB’s additional reconstruction error compared to SVD can be found in Appendix A.
Sparse Orthogonal to Bias (SOB)
When the number of features $p$ exceeds the number of observations $n$, estimating a low-dimensional structure from high-dimensional data can become numerically unstable (Zou et al., 2006). To address this challenge, we introduce a sparse variant of the OB algorithm, referred to as SOB. The SOB imposes an $\ell_{1}$-norm penalty for $\mathbf{U}$ to encourage sparsity and improve numerical stability, in addition to the orthogonality constraints in (5). We define $h$ as the $\ell_{1}$ constraint on $u$. Details of the formulation of SOB can be found in Appendix B. Following derivation by (Witten et al., 2009; Aliverti et al., 2021), Algorithm 1 illustrates the key steps to implement the SOB algorithm, where $\eta$ represents the minimum change to terminate the iterative optimization process.
Note that with the additional regularization constraints in this case, the solution favors sparsity while satisfying the orthogonal constraint. Therefore, SOB also achieves at counterfactual fairness under SCM framework with Theorem 3.3.
Algorithm 1 Sparse Orthogonal to Bias (SOB)
@@
21:Finally calculate the attribute matrix $\widetilde{\mathbf{A}}=\mathbf{\hat{S}}\mathbf{\hat{U}}^{\top}$.
@@
4. Experiments
In this section, we present the results of our experiments conducted on both synthetic and real data sets. In Section 4.2 and 4.3, we compare the proposed OB with several existing methods: Machine Learning (ML), a logistic regression baseline which uses all attributes whether sensitive or not; Fairness through Unawareness (FTU), which simply fits a logistic model with $\mathbf{A}$, excluding the sensitive variables from the model; Equalized Odds (EO), a post-processing algorithm chosen to balance false positives and false negatives while minimizing the expected loss proposed by (Wang et al., 2019); FairLearning Algorithm (FL), which achieves counterfactual fairness by sampling unobserved parts of the graphical causal model using Markov chain Monte Carlo methods (Kusner et al., 2017); Affirmative Action (AA), an post processing algorithm that produces fair equalized odds (EO) and affirmative action (AA) predictors by positing a causal model and considering counterfactual decisions (Wang et al., 2019); Fair Learning through Data pre-processing (FLAP) by (Haoyu Chen and Ghosh, 2023).
[…]
4.1. Evaluation metrics
We assess the accuracy of the decisions concerning the ground truth with Area Under the Curve (AUC) and Accuracy (ACC). As for the evaluation of counterfactual fairness, we adopt two metrics introduced by (Haoyu Chen and Ghosh, 2023): CF-metrics measures counterfactual fairness by calculating the average change in predicted scores between groups with the most significant difference. Additionally, we incorporate CF Bound to evaluate counterfactual fairness in the absence of non-sensitive conditions. This metric computes the maximum absolute value of the bounds’ average of predicted scores for a sample randomly selected from the set. Due to computational challenges, we only evaluate CF-Bound for the simulated data set. For a comprehensive comparison of the methods, we additionally incorporate two observational fairness metrics: EO Fairness, as defined by (Wang et al., 2019), and AA Fairness, proposed in (Wang et al., 2019). In Tables 2, 3 and 1, we highlight the best-performing method in bold and underline the second-best for each metric used.
4.2. Synthetic data
@@
We first apply our methods to a synthetic loan data set example which is a modification from (Haoyu Chen and Ghosh, 2023). Using synthetic data allows us to repeat the random data generation process and provide the average results. It also enables us to observe how fair learning models respond to changing effects resulting from different levels of unfair treatment among different groups. The presented example illustrates a scenario in which a bank evaluates loan applications based on the applicant’s education level ($E$) and annual income ($I$), determining approval ($Y=1$) or rejection ($Y=0$). The population comprises three possible race groups: $B=\{0,1,2\}$. Similar to (Haoyu Chen and Ghosh, 2023), we generate $B$ according to $B=\mathbbm{1}\{U_{B}<0.76\}+\mathbbm{1}\{U_{B}>0.92\}$, where $U_{B}\sim\text{Uniform}(0,1)$. Let $U_{E}$ and $U_{I}$ be two standard normal random variables with mean $\mu_{E}=\lambda_{E0}+\mathbbm{1}\{B=1\}\lambda_{E1}+\mathbbm{1}\{B=2\}\lambda_{E2}$ and $\mu_{I}=\log(\lambda_{A0}+\mathbbm{1}\{B=1\}\lambda_{A1}+\mathbbm{1}\{B=2\}\lambda_{A2})$, respectively. Then the education year and annual income for each race group follows the following distribution:
(8) $E$ $=\max\{0,U_{E}\},$
@@
(9) $Y=$ $\mathbbm{1}\{U_{Y}<\operatorname{expit}(\beta_{0}+\beta_{1}\mathbbm{1}\{B=1\}+\beta_{2}\mathbbm{1}\{B=2\}+\beta_{E}E+\beta_{I}I)\},$
@@
Figure 3. Synthetic loan data. The x-axis shows different levels of the effect of the sensitive variable on education level, $\beta_{E}$, as defined in (9). With 10 repeated experiments, the lines represent the averaged results for each method, and the bars represent their standard deviation. The red line represents the results of the proposed method OB.
Table 1. Performance comparison of our method and other existing methods with synthetic loan data
Metrics Baselines Compared Methods Ours

### v2

@@ 3 Experiments
Table 3: Performance comparison on Adult data [30]
Metrics Baselines Compared Methods Ours
[…]
Table 4: Performance comparison on Synthetic Loan data
Metrics Baselines Compared Methods Ours
[…]
EO Fairness 0.5469 0 0.0156 0 0.0336 0.0321 0.0156 0.0301 0.0180 0 0
Results with discrete sensitive variables
We compare our method with the following baselines on datasets that include discrete sensitive variables: Machine Learning (ML), a straightforward logistic regression model that uses all variables, regardless of their sensitivity; Fairness through Unawareness (FTU), which employs a logistic model with $\mathbf{A}$, specifically excluding sensitive variables; FairLearning Algorithm (FL), which attains counterfactual fairness by sampling unobserved portions of the causal model using Markov chain Monte Carlo methods [7]; EO-fair and AA-fair Predictors (EO and AA) by a post-processing algorithm that generate optimal predictors adhering to Equalized Opportunities (EO) and Affirmative Action (AA) [9]; and Fair Learning through Data Pre-processing (FLAP) by [10], which modifies data prior to model training to enhance fairness. Logistic regression predictors are adopted for all compared models with classification tasks.
[…]
As shown in Table 2 and Table 3, the accuracy of OB is consistently among the highest, often ranking first or second. This supports our claim that the proposed OB algorithm effectively preserves most of the information in the non-sensitive data. Additionally, its CF-metric is comparable to those of FLAP and FL, which are also designed for counterfactual fairness with discrete sensitive variables, indicating that OB also maintains counterfactual fairness. Furthermore, OB achieves relatively low EO and AA Fairness metrics compared to FLAP and FL. Although our is primarily designed to address counterfactual fairness, it also tends to induce a low EO metric, as explained in Appendix E.
In summary, across all discrete datasets, our approach consistently demonstrates its effectiveness in achieving a better overall balance between accuracy and counterfactual fairness, even when the data does not fully meet our theoretical assumptions.
Sensitivity analysis
Table 5: Sensitivity analysis on the COMPAS dataset
Metrics $\text{OB}_{1}$ $\text{OB}_{2}$
$k$ = 3 $k$ = 4 $k$ = 5 $k$ = 3 $k$ = 4 $k$ = 5
ACC 0.5409 0.5507 0.5666 0.5420 0.5516 0.5674
CF-metric 0.0120 0.0157 0.0060 0.0129 0.0162 0.0065
EO Fairness 0 0 0 0 0 0
AA Fairness 0.0006 0.0003 1.6e-16 0.0006 0.0003 1.1e-16
We also investigate the sensitivity of our OB method to the hyperparameter rank $k$. Note that rank $k$ plays a key role in information preservation and data approximation. For SOB, additional hyperparameters, such as $\eta$—used in Algorithm 1 to set the convergence threshold—are involved in solving the optimization problem presented in (6). However, these do not significantly impact the outcomes.
We utilize the COMPAS dataset for sensitivity analysis, which contains 5 features in addition to the sensitive variable. The limited number of feature set makes it easier to illustrate the different effects of varying $k$ by restricting $k$ to a maximum of 5. As shown in Table 5, there is an empirical trend of improved accuracy (ACC), similar to what is observed in Principal Component Analysis (PCA) [23]. Additionally, the counterfactual fairness metrics (CF-metric and AA Fairness) also improve with a higher $k$.
4 Conclusion
In conclusion, this paper demonstrates that achieving counterfactual fairness is feasible by ensuring the uncorrelation between non-sensitive and sensitive variables under certain conditions. Building on this insight, we present the Orthogonal to Bias (OB) algorithm, a novel approach to addressing fairness challenges in machine learning models. OB achieves counterfactual fairness by decorrelating data from sensitive variables under mild conditions. The resulting data pre-processing algorithm effectively removes bias in predictions while making minimal changes to the original data. Importantly, OB is model-agnostic, ensuring its adaptability to a variety of machine learning models. Through comprehensive evaluations on simulated and real-world datasets, we demonstrate that OB strikes a great balance between fairness and accuracy, outperforming compared methods and offering a promising solution to the complex issue of bias in machine learning. Despite the advantages of the OB algorithm, it exhibits certain limitations. As an offline, pre-processing approach, it may not integrate as seamlessly with dynamic or real-time systems compared to in-processing methods. Additionally, the theoretical guarantees of OB depend on jointly normal assumption which may not hold in some practical scenarios.
Appendix A Analyzing the Joint Normality of Linear Combinations in SCM
This section demonstrates how the Structural Causal Model (SCM) depicted in Figure 2(a) aligns with Assumption 2.2 under certain mild conditions. For simplification, we assume both $A$ and $B$ are one-dimensional variables.
Consider a scenario where the underlying exogenous variables $U_{A}$ and $U_{B}$ are distributed as follows:
$U_{A}\sim N(\mu_{A},\sigma_{A}^{2}),$
$U_{B}\sim N(\mu_{B},\sigma_{B}^{2}),$
and $U_{B}$ is independent of $U_{A}$.
Given their normal distributions and independence, $[U_{A},U_{B}]$ form a jointly normal distribution. This property of normal distributions ensures that any linear combination of $U_{A}$ and $U_{B}$ will also be jointly normal. We define $A$ and $B$ through the following linear combinations:
$A$ $=aU_{A}+bU_{B}+c,$
$B$ $=dU_{B}+e,$
where $a$, $b$, $c$, $d$, and $e$ are constants. This formulation ensures that $[A,B]$ is also jointly normal, as long as $A$ and $B$ are defined as linear combinations of $U_{A}$ and $U_{B}$ as described. The synthetic data used for continuous sensitive variables, described in Appendix D.3, follows a similar structure, satisfying Assumption 2.2.
Appendix B Closed-form OB Solution derivation
$\arg\min_{S,U}\left\{\frac{1}{n}\sum_{i=1}^{n}\lVert{a_{i}-\sum_{j=1}^{k}s_{ij}u_{j}^{T}}\rVert^{2}+\frac{2}{n}\lambda_{1}\sum_{i=1}^{n}s_{i1}b_{i}\right\}.$ (7)
Solving it finds a stationary point of
$s_{i1}=a_{i}u_{1}^{T}-\lambda_{1}b_{i}.$ (8)
So the optimal score for the $i$-th subject is obtained by projecting the observed data onto the first basis and then subtracting $\lambda_{1}b$. The constraint does not involve the orthonormal basis $u_{1}$, hence the solution of (7) for $u_{1}$ is equivalent to the unconstrained scenario. A standard result of linear algebra states that the optimal $u_{1}$ for (7) without constraints equivalent to the first right singular vector of $\mathbf{A}$, or equivalently to the first eigenvector of the matrix $\mathbf{A}^{\top}\mathbf{A}$ [26]. Plugging in the solution for $u_{1}$ and setting the derivative with respect to $\lambda_{1}$ equal to 0 leads to
$\sum_{i=1}^{n}(a_{i}u_{1}^{T}-\lambda_{1}a_{i})^{T}b_{i}=0.$ (9)
Therefore,
$\quad\lambda_{1}=\frac{\sum_{i=1}^{n}a_{i}u_{1}^{T}b_{i}}{\sum_{i=1}^{n}b_{i}^{2}}=\frac{\langle Au_{1}^{T},b\rangle}{\langle b,b\rangle},$ (10)
which states $\lambda$ is is a least squares estimate of $\mathbf{A}u^{\top}_{1}$ over $b$.
$s_{ij}=a_{i}u_{j}^{T}-\lambda_{j}b_{i},$ (11)
since $u_{i}^{T}u_{j}=0$ for all $i\neq j$ and $u_{j}^{T}u_{j}=1$ for $j=1,....,k$.
$\lambda_{k}=\frac{\langle Au_{k}^{T},b\rangle}{\langle b,b\rangle}.$ (12)
This results in the closed-form solution in (6). For a more complete proof and discussion of the implications of the solution, we refer to [23]. For example, as noted by [23], an intuitive interpretation of the solution in (11) is that the optimal scores for the $j$-th dimension are obtained by projecting the original data over the $j$-th basis and then subtracting $j$-times the observed value of $b$. Moreover, as the constraints of OB do not involve any vector $u_{j}$, the optimization with respect to the basis can be derived from known results in linear algebra. The optimal value for the vector $u_{j}$, with $j=1,....,k$, is equal to the first k right singular values of $A$, sorted accordingly to the associated singular values [31, 32].
We note the following useful Lemma adapted from [23] that quantifies the additional reconstruction error of $A$ due to using OB compared to SVD is:
Lemma B.1.
Appendix C Formulation of SOB
$\begin{gathered}\underset{S,U}{\arg\min}\left\lVert A-SU^{T}\right\rVert_{F}^{2}\\ \text{ subject to }\left\|u_{j}\right\|_{2}\leq 1,\left\|u_{j}\right\|_{1}\leq t,\left\|s_{j}\right\|_{2}\leq 1,s_{j}^{T}s_{l}=0,s_{j}^{T}B=0,\end{gathered}$ (13)
for $j=1,\dots,k$, and $l\neq j$. The detailed iterative approach to solving this problem is outlined and explained in [23]. The main idea is that although the minimization problem is not jointly convex in $s$ and $u$, it can be addressed iteratively. When $s$ is fixed, the minimization step is equivalent to a sparse matrix decomposition with constraints on the right singular vectors of $A$. On the other hand, when $u$ is fixed, the solution for $s$ is obtained by rearranging the constraints and solving a univariate optimization problem. This iterative process ensures orthogonality among the vectors $s_{j}$.
@@ Appendix C Formulation of SOB
21:  Calculate the attribute matrix $\widetilde{\mathbf{A}}=\mathbf{\hat{S}}\mathbf{\hat{U}}^{\top}$.
@@
Algorithm 1 Sparse Orthogonal to Bias (SOB)
Appendix D Experiment Details
The experiments were conducted in a Jupyter Notebook environment with 16 GB RAM and a 16 GB T4 GPU. Each experiment for the continuous case (Crime and Synthetic) was repeated 10 times to calculate the mean and standard deviation. For regression tasks (continuous $Y$), we utilized a uniform four-layer NN and MSE loss, trained for 2000 epochs with a batch size of 256. Adv and ML have an input size of $q+1$, while FTU, OB, and SOB have an input size of $q$. For classification tasks, we employed logistic regression. The hyperparameter $K$ for our method was set to 4 for the Crime dataset and 20 for the Synthetic insurance price dataset.
D.1 Evaluation Metrics
CF-Metric in Continuous Case
The Counterfactual Fairness (CF) metric for the continuous case, adopted from [10], quantifies the fairness of decisions made by a machine learning model. This metric assesses the difference in predicted outcomes under hypothetical scenarios where a sensitive attribute $\mathbf{B}$ of an individual might differ, while all other attributes remain constant. It is defined as:
$CF=\frac{1}{m_{\text{test}}}\sum_{i=1}^{m_{\text{test}}}\mathbb{E}_{(\mathbf{A}^{\prime}_{i},\textbf{b}^{\prime})\sim C(i)}\left[\left|\Delta(f_{\hat{Y}}(\theta(\mathbf{A}_{i},\mathbf{b}_{i})),f_{\hat{Y}}(\theta(\mathbf{A}^{\prime}_{i},\mathbf{b}^{\prime}_{i})))\right|\right],$
where $m_{\text{test}}$ is the number of test instances, and $C(i)$ denotes the set of counterfactual samples for the $i$-th test instance. These samples are generated based on an adversarial inference process. $\Delta$ represents a cost function comparing two predictions, and $\theta$ denotes the transformation model applied to the inputs. For Adv [12], it’s referring to the VAE. And for our method, it’s the results of OB process. This metric evaluates the average absolute difference between the predicted outcomes under actual and counterfactual attributes, aiming to assess how decisions might vary with changes in the sensitive attribute $B$, while keeping other variables constant. This measure seeks to ensure decisions remain invariant under hypothetical alterations of sensitive attributes, thus quantifying the model’s counterfactual fairness.
CF-Metric in Discrete Case
The CF-metric for the discrete case is designed to measure how decisions made by a machine learning model might differ if a sensitive attribute $B$ were altered, assuming all other attributes $A$ remain unchanged. It is defined as:
$CF=\max_{r,h\in\mathcal{B}}\left(\frac{1}{m_{\text{test}}}\sum_{i=1}^{m_{\text{test}}}\left|f_{\hat{Y}}(\theta(r,\hat{a}_{M}^{D}(r,b_{i},a_{i})))-f_{\hat{Y}}(\theta(b,\hat{a}_{M}^{D}(h,b_{i},a_{i})))\right|\right),$
where $m_{\text{test}}$ is the number of test instances, and $\mathcal{B}$ represents the set of sensitive groups under consideration. $\hat{a}_{M}^{D}(h,b^{*},a^{*})$ is the mapping function to compute non-sensitive attributes assuming that the individual belongs to a different sensitive group $h$ [10]. The functions $f_{\hat{Y}}(\theta(r,\cdot))$ and $f_{\hat{Y}}(\theta(h,\cdot))$ evaluate decision outcomes for real and hypothetical scenarios, respectively. This metric aims for a value of zero, indicating perfect counterfactual fairness across all considered groups.
For AA and EO metric definitions, we refer to [9] for details.
D.2 Real-world Datasets
Adult
In the Adult Income dataset [30], we aim to predict whether an individual’s income exceeds \$50,000, considering features such as sex, race, age, work class, education, occupation, marital status, capital gain, and loss. The sensitive variables are sex and race. The training set consists of 32,561 samples, and the test set comprises 16,281 samples.
Due to the large sample size of Adult dataset, we employ SOB. As illustrated in Table 3, similar to OB’s performance in synthetic data, the accuracy is comparatively high compared to all other tested fair learning approaches. Notably, the accuracy is even higher than the vanilla ML model, which utilizes both sensitive and non-sensitive variables and generates unfair results. As noted by [23, 31], the additional regulation with SOB may contribute to high out-of-sample prediction performances. Moreover, its CF-metric is comparable to that of FLAP and FL and is much lower than baselines, implying counterfactual fairness attributed to OB. Additionally, it achieves both low EO and AA Fairness metrics.
COMPAS
The COMPAS recidivism data [28] includes demographic information such as sex, age, race, and record data (prior counts, juvenile felonies counts, and juvenile misdemeanors counts) for over 10,000 criminal defendants in Broward County, Florida. The goal is to predict whether they will re-offend in the next two years.
Crime
Crime dataset [29] with the ratio of an ethnic group per population as sensitive variable, and violent crimes per population as Y.
D.3 Simulated Datasets
D.3.1 Synthetic Insurance Price Dataset
We simulate two scenarios for continuous sensitive variable $B$ and decision variable $Y$ following Figure 2(a):
$U\sim\mathcal{N}\left(\begin{bmatrix}0\\ 0.5\\ 1\\ 1.5\\ 2\\ \end{bmatrix},\begin{bmatrix}1&0&0&0&0\\ 0&4&0&0&0\\ 0&0&2&0&0\\ 0&0&0&3&0\\ 0&0&0&0&2\\ \end{bmatrix}\right)$
The non-sensitive variables $A_{1},A_{2},A_{3},A_{4}$ are defined as follows:
$A_{1}\sim\mathcal{N}(7+0.1\cdot B+U_{1}+U_{2}+U_{3},1),$
$A_{2}\sim\mathcal{N}(80+B+U_{2},10),$
$A_{3}\sim\mathcal{N}(200+5\cdot B+5\cdot U_{3},20),$
$A_{4}\sim\mathcal{N}(10^{4}+5\cdot B+U_{4}+U_{5},1000),$
where $B$ is defined by the vector:
$B\sim\mathcal{N}(\begin{bmatrix}45,5\end{bmatrix}).$
The response variable $Y$ is modeled as:
$Y\sim\mathcal{N}(2\cdot(7\cdot B+20\cdot\sum_{j}A_{j}),0.1).$
D.3.2 Synthetic Loan Decision Dataset
@@
We apply our methods to a synthetic loan dataset example which is a modification from [10]. Using synthetic data allows us to repeat the random data generation process and provide the average results. It also enables us to observe how fair learning models respond to changing effects resulting from different levels of unfair treatment among different groups. The presented example illustrates a scenario in which a bank evaluates loan applications based on the applicant’s education level ($E$) and annual income ($I$), determining approval ($Y=1$) or rejection ($Y=0$). The population comprises three possible race groups: $B=\{0,1,2\}$. Similar to [10], we generate $B$ according to $B=\mathbbm{1}\{U_{B}<0.76\}+\mathbbm{1}\{U_{B}>0.92\}$, where $U_{B}\sim\text{Uniform}(0,1)$. Let $U_{E}$ and $U_{I}$ be two standard normal random variables with mean $\mu_{E}=\lambda_{E0}+\mathbbm{1}\{B=1\}\lambda_{E1}+\mathbbm{1}\{B=2\}\lambda_{E2}$ and $\mu_{I}=\log(\lambda_{A0}+\mathbbm{1}\{B=1\}\lambda_{A1}+\mathbbm{1}\{B=2\}\lambda_{A2})$, respectively. Then the education year and annual income for each race group follows the following distribution:
$E$ $=\max\{0,U_{E}\},$ (14)
@@
$Y=$ $\mathbbm{1}\{U_{Y}<\operatorname{expit}(\beta_{0}+\beta_{1}\mathbbm{1}\{B=1\}+\beta_{2}\mathbbm{1}\{B=2\}+\beta_{E}E+\beta_{I}I)\},$ (15)
@@
It is important to note that the sensitive variable $B$ is categorical, and the data generating process does not exactly conform to Assumption 2.2. As evidenced in Table 4, despite the deviation from the assumptions in the tested synthetic dataset, our method consistently showcases comparatively high AUC and ACC compared to most methods. Notably, its accuracy outperforms FL and FLAP, two other counterfactually fair methods. Moreover, our method achieves low CF-metric, akin to FL and FLAP, indicating a high degree of counterfactual fairness. This desirable characteristic can be attributed to the property of OB, which minimally modifies in the non-sensitive data while ensuring counterfactual fairness. Furthermore, in terms of observational fairness metrics, OB exhibits an overall better performance compared to FL and FLAP with lower EO and AA Fairness metrics.
Appendix E Explanation for EO Fairness of OB and SOB